OpenStudy (anonymous):

Eliminate the parameter of x =1/2 t, y=((3t)^3) -1

OpenStudy (loser66):

Your avatar is so cute!!

OpenStudy (loser66):

\[x =\dfrac{t}{2}\rightarrow t=??\]

OpenStudy (anonymous):

Thanks! And are you asking for clarification of the problem or is that the first step to solving it?

OpenStudy (loser66):

That is the first step

OpenStudy (anonymous):

But x is equal to 1/2 not t/2?

OpenStudy (loser66):

1/2 t is 1/2*t = t/2 hun

OpenStudy (anonymous):

K but my equation looks like this: http://imgur.com/NM4EDJc I think I wrote it out wrong...

OpenStudy (loser66):

You are good, just find t and then plug t into y Ok, I do it for you. \[x =\frac{t}{2}\rightarrow t = 2x\] \[y =(3t)^3 -1\] replace t =2x above into y \[y =(3(2x))^3 -1 =(6x)^3 -1=216x^3 -1\] no more t, right? That's the process to "Eliminate t"

OpenStudy (anonymous):

Sorry, i had to step out to talk to my teacher. That makes a lot of sense (and now i feel really stupid) whats the next step?

OpenStudy (loser66):

the next step is "submit"

OpenStudy (anonymous):

I don't have your answer exactly but I have y = 6x^3 - 1? 6 is the cube root of 216 so would that work?

OpenStudy (anonymous):

Thanks so much for all your help! <3